# SigmaDock: Untwisting Molecular Docking With Fragment-Based $\operatorname{SE}(3)$ Diffusion

### v1

@@ Abstract
Determining the binding pose of a ligand to a protein, known as molecular docking, is a fundamental task in drug discovery. Generative approaches promise faster, improved, and more diverse pose sampling than physics-based methods, but are often hindered by chemically implausible outputs, poor generalisability, and high computational cost. To address these challenges, we introduce a novel fragmentation scheme, leveraging inductive biases from structural chemistry, to decompose ligands into rigid-body fragments. Building on this decomposition, we present SigmaDock, an $\operatorname{SE}(3)$ Riemannian diffusion model that generates poses by learning to reassemble these rigid bodies within the binding pocket. By operating at the level of fragments in $\operatorname{SE}(3)$, SigmaDock exploits well-established geometric priors while avoiding overly complex diffusion processes and unstable training dynamics. Experimentally, we show SigmaDock achieves state-of-the-art performance, reaching Top-1 success rates (RMSD <2<2 & PB-valid) above 79.9% on the PoseBusters set, compared to 12.7-30.8% reported by recent deep learning approaches, whilst demonstrating consistent generalisation to unseen proteins. SigmaDock is the first deep learning approach to surpass classical physics-based docking under the PB train-test split, marking a significant leap forward in the reliability and feasibility of deep learning for molecular modelling.
@@ 1 Introduction
The biological function of a protein is determined primarily by its 3D structure and the interactions it mediates. Thus, a central goal of drug discovery is to design small-molecule ligands that bind to a target protein and modulate its function to achieve a therapeutic effect. Since a change in protein function (activation or deactivation) is highly correlated with the bound pose of the binding ligand, a consequence of energetically favourable protein-ligand interactions, the ability predict these structural conformations, which is the primary aim of molecular docking, is essential for reliable and accelerated drug discovery.
Deep learning approaches for molecular docking, in particular, diffusion-based methods, have been recently touted as providing superior accuracy over traditional, industry-standard physics-based methods (Halgren et al., 2004; Morris & Lim-Wilby, 2008; Trott & Olson, 2010). However, the need for such claims to be further qualified has been highlighted by Harris et al. (2023); Buttenschoen et al. (2024), who demonstrated that solely focusing on metrics, such as Root Mean Squared Deviation (RMSD) between the bound and predicted ligand poses, can obfuscate the actual predictive ability of deep learning-based docking tools. For instance, Buttenschoen et al. (2024) showed that when controlling for the chemical plausibility of generated samples, deep learning approaches performed far worse than traditional docking methods. While notable progress on this front has been made with the advent of co-folding models (Abramson et al., 2024; Boitreaud et al., 2024; Wohlwend et al., 2024) such as AlphaFold3 (AF3), these models still have key limitations. Firstly, they require massive quantities of data and compute to train, hindering the ability of other researchers to contribute and improve such models. Secondly, as co-folding models jointly model the structure of proteins and ligands, they suffer from slow inference, which makes their practical applicability to drug discovery computationally prohibitive; especially for virtual screening, where it is often required to query millions of protein-ligand pairs.
@@
Empirically, we demonstrate SigmaDock surpasses prior deep learning approaches [^1] Under fair comparison with models trained on the PoseBusters train-test split and traditional physics-based docking, achieving state-of-the-art performance on the challenging PoseBusters set (Buttenschoen et al., 2024), and generalising to unseen proteins (Figure 4). We highlight that, with a fraction of the training data, training/sampling time, and lower test-train leakage, we reach AF3-level performance. With this, we wish to state our main contribution as the careful and rigorous design of a well-characterised diffusion process, and a detailed construction of structural inductive biases which help learn simpler functions for the task of molecular docking.
@@ 2.1 Notation
2.2 Structurally-Aware Fragmentation for SE(3) Diffusion
@@ 2.2.1 The Conformational Manifold
SigmaDock makes use of well-established thermodynamic priors in structural chemistry. Namely, it leverages the fact that the conformational space $\mathbf{x}_{c}$ defining the local geometry of a ligand (in a vacuum) with known topological structure $\mathcal{G}_{\text{ligand}}^{\text{2D}}$ follows a Boltzmann distribution with probability mass concentrated about a manifold $\mathcal{M}_{c}$ with holonomic constraints [^4] Holonomic constraints restrict the configuration space (position) of a body. of the form:
@@
Empirically, we find that the RMSDs between experimentally bound poses $\mathbf{x}_{b}$ and their corresponding aligned conformers $\mathbf{x}_{b}^{\prime}=\Psi(p^{\star},R^{\star},\bm{\phi}^{\star})$ are substantially below both (i) standard error rates reported by docking baselines (Buttenschoen et al., 2024; Harris et al., 2023), and (ii) the commonly used success threshold of 2Å. This provides sufficient support to claim that the variability in bond lengths and bond angles subsumed in $\mathcal{M}_{c}$ can be generally ignored in the task protein-ligand docking. Crucially, this allows us to treat bound states as being approximately contained in the set of structures reachable by torsions and $\operatorname{SE}(3)$ transforms on conformers drawn from $\pi_{\mathcal{M}_{c}}$: for any $\mathbf{x}_{c}\in\mathcal{M}_{c}$ and $\mathbf{x}_{b}\in\mathcal{M}_{b}$, we can align $\mathbf{x}_{c}$ to $\mathbf{x}_{b}$ with negligible error such that $\text{RMSD}(\mathbf{x}_{b},\mathbf{x}_{b}^{\prime})\ll 2$Å. See Figure 2b for an aligned example and Appendix D.3 for additional results and empirical analysis. This empirical inclusion is paramount to our approach as it justifies assembling our stationary distribution from fragments sampled from $\mathcal{M}_{c}$, without falling out of distribution. Throughout the remainder of the paper, we will absorb the alignment into the notation by writing $\mathbf{x}\leftarrow\mathbf{x}_{b}\leftarrow\mathbf{x}_{b}^{\prime}$.
@@ 2.2.2 Challenges of Torsional Models & Motivation for Our Method
For standard molecular topologies, torsional models define nonlinear mappings from torsion angles to Cartesian coordinates, producing highly entangled, non-product induced measures. In contrast, disjoint rigid fragments yield a factorised product of Haar measures on $\mathrm{SE}(3)^{m}$.
Consequently, we argue that diffusing a molecule in fragment space $\operatorname{SE}(3)^{m}$ offers a simpler learning task than diffusing in torsion space $\mathbb{T}^{k}\times\operatorname{SE}(3)$. Intuitively, local changes in torsional angles often produce non-local Cartesian displacements, creating strong geometric coupling along torsional chains: a change in a single torsion can substantially displace remote atoms. Therefore, transition kernels that perturb torsions independently become geometrically entangled once mapped to Cartesian coordinates (where the model observes the data) via the induced measure, breaking the product structure and leading to ill-conditioned and often degenerate dynamics during training and sampling.
@@ 2.2.3 Irreducible Fragmentation & Soft Geometric Constraints on $\operatorname{SE}(3)$
The naive choice to define our fragments is to break the molecular graph obtained from $\pi_{\mathcal{M}_{c}}$ at the torsional bonds, producing a set $\{\mathcal{G}_{F_{i}}\}_{i=1}^{\hat{m}}$ of torsion-free rigid-body fragments with global coordinates parametrised by $\operatorname{SE}(3)^{\hat{m}}$. This approach yields a set of $\hat{m}=(k+1)$ fragments with a total of $6\hat{m}$ DoFs [^5] Assuming $|\mathcal{G}_{F}|>1$, each fragment is defined by its $\mathbb{\mathbb{T}}(3)\cong\mathbb{R}^{3}$ and $SO(3)$ parametrisation.. In contrast, we note torsional models have $(k+6)$ DoFs ($k$ torsional bonds in $\mathcal{S}^{1}$ and 6 for rigid body SE(3)). Thus, the natural question arises: How can we reduce the DoFs of the system and in turn abstract the problem in a general form? In SigmaDock we tackle this problem by creating a simple yet effective molecular fragmentation reduction (FR3D) that recursively merges adjacent fragments from $\hat{m}=(k+1)$ down to $m$ (Figure 3). Instead of biasing the fragmentation order, FR3D performs a stochastic search, starting from the $\hat{m}$ fragments and branching through candidate neighbours until it reaches an irreducible set of size $m$. Hence, FR3D not only reduces the learnable DoFs but also provides a promising stream of data augmentation. Since merging is only possible in molecular graphs where two or more consecutive torsional bonds are linked, there are topologies that are irreducible, thus we are still upper bound in the number of fragments: $0\leq m\leq k+1$. However, we find FR3D reduces the average number of fragments by $\approx 66\%$.
During fragmentation, we retain torsional bond length and angle information by introducing dummy atoms at either side of the bond. Hence, $|\mathcal{G}_{\text{ligand}}|\leq\sum_{i=1}^{m}|\mathcal{G}_{F_{i}}|$. Importantly, we only retain free dummy atoms in $\mathcal{G}_{F_{i}}$ and otherwise prune dummies which are over-constrained. Here, we label a dummy as over-constrained whenever FR3D merges the torsional bond it belongs with a neighbouring fragment, as it naturally over-defines a dihedral angle (Figure 3b). Removing over-constrained atoms is fundamental since an immutable dihedral sampled from $\pi_{\mathcal{M}_{c}}$ would violate the free torsional requirements outlined in Section 2.2.1, forcing our generator $\pi_{\mathcal{M}_{c}}$ to yield structures that do not strictly overlap with the bound manifold under optimal alignment: $\Psi\cdot\pi_{\mathcal{M}_{c}}(\cdot)\neq\pi_{\mathcal{M}_{b}}(\cdot)$. We refer the reader to Algorithm 1 in Appendix D.4 for an overview of FR3D and further analysis.
@@ 2.4 Architecture
Moreover, we note that a critical issue for the design of our architecture is that the parametrisation of the global coordinates $\mathbf{x}_{F}$ in terms of $(p,R)\in\operatorname{SE}(3)$ is not uniquely defined. This is due to the fact that we do not have a canonical choice for the orientation of the local coordinates $\tilde{\mathbf{x}}_{F}$. For instance, we have the equally valid choices $\tilde{\mathbf{x}}_{F},\tilde{\mathbf{x}}_{F}^{\prime}$ for the local coordinates of $\mathcal{G}_{F}$ if $\tilde{\mathbf{x}}_{F}^{\prime}=R_{0}\cdot\tilde{\mathbf{x}}_{F}$ where $R_{0}\in\operatorname{SO}(3)$. Hence, we can have two different representations of global coordinates $\mathbf{x}_{F}$ from $\mathbf{x}_{F}=(p,R)\cdot\tilde{\mathbf{x}}_{F}=(p,RR_{0}^{-1})\cdot\tilde{\mathbf{x}}_{F}^{\prime}$ depending on the initial choice of orientation. To resolve this issue, we adapt the $\operatorname{SO}(3)$-equivariant prediction head introduced in Jin et al. (2023), based on the Newton-Euler equations from rigid-body mechanics, to which we pass the outputs of our backbone model into. With this choice, Theorem 2 shows that SigmaDock is invariant to the choice of local coordinate axes.
@@ 2.5 Training and Inference
We outline our training setup for SigmaDock in Appendix E. In particular, we discuss how we preprocess our data for fragmentation and training, as well as computational tricks for increasing training throughput. Our sampling procedure is outlined in Appendix F, where we discuss the fact that, due to the reliability of SigmaDock in generating chemically plausible samples, SigmaDock does not require the use of a separately trained confidence model to filter out poor generations. Instead, we propose using the simple and cheap heuristic of evaluating both the (pseudo) binding energy of the generated protein-ligand system, as well as a set of physicochemical checks (such as, bond angles, bond lengths, internal energy) to rank our $N_{\text{seeds}}$ samples for evaluation.

### v2

@@ Abstract
Determining the binding pose of a ligand to a protein, known as molecular docking, is a fundamental task in drug discovery. Generative approaches promise faster, improved, and more diverse pose sampling than physics-based methods, but are often hindered by chemically implausible outputs, poor generalisability, and high computational cost. To address these challenges, we introduce a novel fragmentation scheme, leveraging inductive biases from structural chemistry, to decompose ligands into rigid-body fragments. Building on this decomposition, we present SigmaDock, an $\operatorname{SE}(3)$ Riemannian diffusion model that generates poses by learning to reassemble these rigid bodies within the binding pocket. By operating at the level of fragments in $\operatorname{SE}(3)$, SigmaDock exploits well-established geometric priors while avoiding overly complex diffusion processes and unstable training dynamics. Experimentally, we show SigmaDock achieves state-of-the-art performance, reaching Top-1 success rates (RMSD <2<2 & PB-valid) above 79.9% on the PoseBusters set, compared to 12.7-  32.8% reported by recent deep learning approaches, whilst demonstrating consistent generalisation to unseen proteins. SigmaDock is the first deep learning approach to surpass classical physics-based docking under the PB train-test split, marking a significant leap forward in the reliability and feasibility of deep learning for molecular modelling.
@@ 1 Introduction
The biological function of a protein is determined primarily by its 3D structure and the interactions it mediates. Thus, a central goal of drug discovery is to design small-molecule ligands that bind to a target protein and modulate its function to achieve a therapeutic effect. Since a change in protein function is highly correlated with the bound pose of the binding ligand, a consequence of energetically favourable protein-ligand interactions, the ability predict these structural conformations, which is the primary aim of molecular docking, is essential for reliable and accelerated drug discovery.
Deep learning approaches for molecular docking, in particular, diffusion-based methods, have been recently touted as providing superior accuracy over traditional, industry-standard physics-based methods (Halgren et al., 2004; Morris & Lim-Wilby, 2008; Trott & Olson, 2010). However, the need for such claims to be further qualified has been highlighted by Harris et al. (2023); Buttenschoen et al. (2024), who demonstrated that solely focusing on metrics, such as Root Mean Squared Deviation (RMSD) between the bound and predicted ligand poses, can obfuscate the actual predictive ability of deep learning-based docking tools. For instance, Buttenschoen et al. (2024) showed that when controlling for the chemical plausibility of generated samples, deep learning approaches performed far worse than traditional docking methods. While notable progress on this front has been made with the advent of co-folding models (Abramson et al., 2024; Boitreaud et al., 2024; Wohlwend et al., 2024) such as AlphaFold3 (AF3),   which need not assume the protein to be a rigid structure, these models still have key limitations. Firstly, they require massive quantities of data and compute to train, hindering the ability of other researchers to contribute and improve such models. Secondly, as co-folding models jointly model the structure of proteins and ligands, they suffer from slow inference, which makes their practical applicability to drug discovery computationally prohibitive; especially for   high-throughput virtual screening   (HTVS), where it is often required to query millions of protein-ligand pairs.
@@
 We adopt the standard re-docking protocol in which the receptor is fixed (holo-conformation) and the binding pocket is known. This choice is deliberate for two reasons: first, it is the long-standing setting for benchmarking docking methods, and permits fair ’apples-to-apples’ comparison with classical and recent deep learning baselines; second, it reflects industrial HTVS/lead-optimisation practice, where rigid-receptor docking is computationally tractable at scale. Prior generative methods have struggled in precisely this regime, and our aim is to close that gap first.
Empirically, we demonstrate SigmaDock surpasses prior deep learning approaches [^1] Under fair comparison with models trained on the PoseBusters train-test split. and traditional physics-based docking, achieving state-of-the-art performance on the challenging PoseBusters set (Buttenschoen et al., 2024), and generalising to unseen proteins (Figure 4). We highlight that, with a fraction of the training data, training/sampling time, and lower test-train leakage, we reach AF3-level performance and   substantially outperform previous generative methods on the re-docking task. With this, we wish to state our main contribution as the careful and rigorous design of a well-characterised diffusion process, and a detailed construction of structural inductive biases which help learn simpler functions for the task of molecular docking.   We highlight that our fragment $\operatorname{SE}(3)^{m}$ formulation naturally extends to flexible docking by treating selected side chains as additional fragments, and is also adaptable to co-folding, both of which we leave as future work.
@@ 2.1 Notation
2.2 Structurally-Aware Fragmentation for $\operatorname{SE}(3)$ Diffusion
@@ 2.2.1 The Conformational Manifold
SigmaDock makes use of well-established thermodynamic priors in structural chemistry. Namely, it leverages the fact that the conformational space $\mathbf{x}_{c}$ defining the local geometry of a ligand (in a vacuum) with known topological structure $\mathcal{G}_{\text{ligand}}^{\text{2D}}$ follows a Boltzmann distribution with probability mass concentrated about a manifold $\mathcal{M}_{c}$ with holonomic constraints [^4] Holonomic constraints (  see (Ryckaert et al., 1977)) restrict the configuration space (position) of a body. of the form:
@@
Empirically, we find that the RMSDs between experimentally bound poses $\mathbf{x}_{b}$ and their corresponding aligned conformers $\mathbf{x}_{b}^{\prime}=\Psi(p^{\star},R^{\star},\bm{\phi}^{\star})$ are substantially below both (i) standard error rates reported by docking baselines (Buttenschoen et al., 2024; Harris et al., 2023), and (ii) the commonly used success threshold of 2Å. This provides sufficient support to claim that the variability in bond lengths and bond angles subsumed in $\mathcal{M}_{c}$ can be generally ignored in the task protein-ligand docking. Crucially, this allows us to treat bound states as being approximately contained in the set of structures reachable by torsions and $\operatorname{SE}(3)$ transforms on conformers drawn from $\pi_{\mathcal{M}_{c}}$: for any $\mathbf{x}_{c}\in\mathcal{M}_{c}$ and $\mathbf{x}_{b}\in\mathcal{M}_{b}$, we can align $\mathbf{x}_{c}$ to $\mathbf{x}_{b}$ with negligible error such that $\text{RMSD}(\mathbf{x}_{b},\mathbf{x}_{b}^{\prime})\ll 2$Å. See Figure 2b for an aligned example and Appendix D.3 for additional results and empirical analysis. This empirical inclusion is paramount to our approach as it justifies assembling our stationary distribution from fragments sampled from $\mathcal{M}_{c}$, without falling out of distribution. Throughout the remainder of the paper, we will absorb the alignment into the notation by writing $\mathbf{x}\leftarrow\mathbf{x}_{b}\leftarrow\mathbf{x}_{b}^{\prime}$,   where we use the aligned conformation $\mathbf{x}_{b}^{\prime}$ at the start of forward noising process.
@@ 2.2.2 Challenges of Torsional Models & Motivation for Our Method
For standard molecular topologies, torsional models define nonlinear mappings from torsion angles to Cartesian coordinates, producing highly entangled, non-product induced measures. In contrast, disjoint rigid fragments yield a factorised product of Haar measures on $\operatorname{SE}(3)^{m}$.
Consequently, we argue that diffusing a molecule in fragment space $\operatorname{SE}(3)^{m}$ offers a simpler learning task than diffusing in torsion space $\mathbb{T}^{k}\times\operatorname{SE}(3)$. Intuitively, local changes in torsional angles often produce non-local Cartesian displacements, creating strong geometric coupling along torsional chains: a change in a single torsion can substantially displace remote atoms. Therefore,   independent torsional perturbations become correlated once mapped to Cartesian coordinates (where the model observes the data)   under the induced measure, breaking the product structure.   This leads to an ill-conditioned   learning problem and stiff sampling dynamics.   In contrast, our forward kernel factorises over disjoint $\operatorname{SE}(3)^{m}$ fragments (product); inter-fragment correlations enter only via the learnt score, rather than being induced by the noise, yielding simpler, better-conditioned mappings, and more stable reverse-time integration.
@@ 2.2.3 Irreducible Fragmentation & Soft Geometric Constraints on $\operatorname{SE}(3)$
The naive choice to define our fragments is to break the molecular graph obtained from $\pi_{\mathcal{M}_{c}}$ at the torsional bonds, producing a set $\{\mathcal{G}_{F_{i}}\}_{i=1}^{\hat{m}}$ of torsion-free rigid-body fragments with global coordinates parametrised by $\operatorname{SE}(3)^{\hat{m}}$. This approach yields a set of $\hat{m}=(k+1)$ fragments with a total of $6\hat{m}$ DoFs [^5] Assuming $|\mathcal{G}_{F}|>1$, each fragment is defined by its $\mathbb{\mathbb{T}}(3)\cong\mathbb{R}^{3}$ and $\operatorname{SO}(3)$ parametrisation.. In contrast, we note torsional models have $(k+6)$ DoFs ($k$ torsional bonds in $\mathcal{S}^{1}$ and 6 for rigid body $\operatorname{SE}(3)$). Thus, the natural question arises: How can we reduce the DoFs of the system and in turn abstract the problem in a general form? In SigmaDock we tackle this problem by creating a simple yet effective molecular fragmentation reduction (FR3D) that recursively merges adjacent fragments from $\hat{m}=(k+1)$ down to $m$ (Figure 3).
Instead of biasing the fragmentation order, FR3D performs a stochastic search, starting from the   torsion-free $\hat{m}$ fragments and branching through candidate neighbour   proposing merge actions until reaching an irreducible set of size $m$. Hence, FR3D not only reduces the learnable DoFs   by reducing the number of fragments, but also provides a promising stream for data augmentation. Merging is possible in molecular graphs where two or more consecutive torsional bonds are linked, so there are topologies that are irreducible. Hence, we are upper-bounded in the number of fragments: $1\leq m\leq k+1$.   For a fragment hyper-graph with $m$ fragments (and no loop closures), the effective DoFs concentrate between k+6k\!+\!6 (triangulation-induced lower bound) [^6] The resulting DoFs are lower-bounded because the triangulation scheme imparts soft boundary constraints; it provides a   strong signal for SigmaDock to reduce   $\Delta d_{A,C}$ to 0 as $t\rightarrow 0$. and $6m$ (unconstrained). Under naïve fragmentation ($\hat{m}\!=\!k\!+\!1$), this becomes $k\!+\!6\leq\mathrm{DoF}\leq 6\hat{m}$. As FR3D reduces $m$ (empirically we find $m\!\approx\!\tfrac{2}{3}\hat{m}$), the upper bound shrink in practice. We refer the reader to Appendix D.4 to a more extensive analysis and empirical results.
During fragmentation, we retain torsional bond length and angle information by introducing dummy atoms at either side of the bond. Hence, $|\mathcal{G}_{\text{ligand}}|\leq\sum_{i=1}^{m}|\mathcal{G}_{F_{i}}|$. Importantly, we only retain free dummy atoms in $\mathcal{G}_{F_{i}}$ and otherwise prune dummies which are over-constrained. Here, we label a dummy as over-constrained whenever FR3D merges the torsional bond it belongs with a neighbouring fragment, as it naturally over-defines a dihedral angle (Figure 3b). Removing over-constrained atoms is fundamental since an immutable dihedral sampled from $\pi_{\mathcal{M}_{c}}$ would violate the free torsional requirements outlined in Section 2.2.1, forcing our generator $\pi_{\mathcal{M}_{c}}$ to yield structures that do not strictly overlap with the bound manifold under optimal alignment: $\Psi\cdot\pi_{\mathcal{M}_{c}}(\cdot)\neq\pi_{\mathcal{M}_{b}}(\cdot)$.
We refer the reader to Algorithm 1 in Appendix D.4 for an overview of FR3D and further analysis.
@@ 2.4 Architecture
Moreover, we note that a critical issue for the design of our architecture is that the parametrisation of the global coordinates $\mathbf{x}_{F}$ in terms of $(p,R)\in\operatorname{SE}(3)$ is not uniquely defined. This is due to the fact that we do not have a canonical choice for the orientation of the local coordinates $\tilde{\mathbf{x}}_{F}$. For instance, we have the equally valid choices $\tilde{\mathbf{x}}_{F},\tilde{\mathbf{x}}_{F}^{\prime}$ for the local coordinates of $\mathcal{G}_{F}$ if $\tilde{\mathbf{x}}_{F}^{\prime}=R_{0}\cdot\tilde{\mathbf{x}}_{F}$ where $R_{0}\in\operatorname{SO}(3)$. Hence, we can have two different representations of global coordinates $\mathbf{x}_{F}$ from $\mathbf{x}_{F}=(p,R)\cdot\tilde{\mathbf{x}}_{F}=(p,RR_{0}^{-1})\cdot\tilde{\mathbf{x}}_{F}^{\prime}$ depending on the initial choice of orientation. To resolve this issue, we adapt the $\operatorname{SO}(3)$-equivariant prediction head introduced in Jin et al. (2023), based on the Newton-Euler equations from rigid-body mechanics, to which we pass the outputs of our backbone model into.   Particularly, we predict pseudo-forces for all atoms pertaining to the $m$ fragments and use these as a basis to define our scores in the tangent space of $\operatorname{SE}(3)$ (more details in Appendix G.4). With this choice, Theorem 2 shows that SigmaDock is invariant to the choice of local coordinate axes.
@@ 2.5 Training and Inference
We outline our training setup for SigmaDock in Appendix E. In particular, we discuss how we preprocess our data for fragmentation and training, as well as computational tricks for increasing training throughput.   We detail the definition of the binding pocket in Appendix E.1. Briefly, the pocket includes all residues with any atom within a stochastic cutoff $d_{\text{r}}$ of any ligand atom. Formally, $d_{\text{r}}:=d_{0}+\mathcal{N}(0,\sigma_{r})$, where $d_{0}$ and $\sigma_{r}$ default to 5Å and 1Å respectively during training. Our sampling procedure is outlined in Appendix F, where we discuss the fact that, due to the reliability of SigmaDock in generating chemically plausible samples, SigmaDock does not require the use of a separately trained confidence model to filter out poor generations. Instead, we propose using the simple and cheap heuristic of evaluating both the (pseudo) binding energy of the generated protein-ligand system, as well as a set of physicochemical checks (such as, bond angles, bond lengths, internal energy) to rank our $N_{\text{seeds}}$ samples for evaluation.